Bicycle A bicycle with 40.20 cm diameter wheels is traveling at 28.00 km/hr. At what angular...

Bicycle

A bicycle with 40.20 cm diameter wheels is traveling at 28.00 km/hr. At what angular speed do the wheels turn?

part B: How long do they take to turn once around?
Answer #1

here,
v = 28 km/hr * 18/5 = 7.78 m/s
r = 40.20/2 = 20.1 cm = 0.201 m

as
angular speed = linear speed/radius
w = v/r
w = 7.78/0.201
w = 38.706 rad/s

Now

T = 2*pi*r/v
T = (2*3.14*0.201)/7.78
T = 0.162 s